Exhibit 4.1 The balance sheet and income statement shown below are for Koski Inc. Note that the firm has no amortization charges, it does not lease any assets, none of its debt must be retired during the next 5 years, and the notes payable will be rolled over Balance Sheet (Millions of $) Assets Cash and securities Accounts receivable Inventories Total current assets Net plant and equipment Total assets Liabilities and Equity 2016 $2,145 8,970 12,480 $23,595 $15,405 $39,000 Accounts payable Accruals Notes payable Total current liabilities $7,410 4,290 5,460 $17,160 Long-term bonds Total liabilities Common stock Retained earnings Total common equity Total liabilities and equity $7,800 $24,960 $5,460 8,580 $14,040 $39,000Refer to

Income Statement (Millions of $) 2016 Net sales Operating costs except depreciation Depreciation Earnings before interest and taxes (EBIT) Less interest Earnings before taxes (EBT) Taxes Net income $58,500 54,698 1,024 $2,779 829 $1,950 683 $1,268 Other data: Shares outstanding (millions) Common dividends (millions of $) Int rate on notes payable & L-T bonds Federal plus state income tax rate Year-end stock pricee 500.00 $443.63 62 % 35% $30.42 Refer to Exhibit 4.1. What is the firms return on invested capital? a. 6.62% b. 4.35% C. 5.05% d. 4.17% e. 4.12% O

Exhibit 4.1. What is the firm's return on invested capital? PLEASE EXPLAIN!
Answer #1

ANSWER is 6.62%

ROIC = NOPAT/Invested capital

NOPAT = EBIT*(1-Tax rate)

= 2779*(1-683/1950)

= 1805.64

Invested capital = Long term debt + Equity + Notes payable

= 7800+14040+5460

=27300

ROIC= 1805.64/ 27300

= 6.62%